OpenStudy (anonymous):

@ikatouni @ipwbunnies Samuel hit the baseball from 2 feet off the ground at an initial vertical velocity of 30 feet per second. How long will it take the ball to hit the ground? A. Write the formula. B. Solve the problem. C. State your answer. Round your answer to the nearest hundredth of a second.

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

umm the equation for height is h= -16t^2+vt+s

OpenStudy (anonymous):

v is velocity s is initial height

OpenStudy (anonymous):

now when ball hits the ground height is 0

OpenStudy (anonymous):

so the equation will be -16t^2+30t+2 = 0

OpenStudy (anonymous):

from this you can find time

OpenStudy (anonymous):

now divide this equation by -2

OpenStudy (anonymous):

you get 8t^2-15t-1=0

OpenStudy (anonymous):

So it would take 0 seconds?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

you have to find T here do you remember yesterday's formula

OpenStudy (anonymous):

wait so I have to factor it, then I get my answer?

OpenStudy (anonymous):

yes but here you have to use \[-b \pm \sqrt{b^2-4c}/2a\]

OpenStudy (anonymous):

Quadratic formula

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so find t ,that will be ur answer